OpenStudy (anonymous):

(14x^3-21x^2+28x)divided by (7x)

OpenStudy (campbell_st):

well you can treat is as 3 separate things \[\frac{14x^3}{7x} -\frac{21x^2}{7x} + \frac{28x}{7x}\] the rule for dividing the same base \[\frac{x^a}{x^b} = x^{a - b}\] subtract the powers... hope this helps.

OpenStudy (anonymous):

it said Find the quotient.

OpenStudy (campbell_st):

well this will give you the quotient... here is the 1st term in the quotient... \[\frac{14x^3}{7x} = 2x^2\]